A lottery has a $ 2 000 000 first prize, a $ 250 000 second prize, and a three $5000 third prizes. A total of 1 000 000 tickets are sold.
a) What is the probability of winning a prize in this lottery?
b) If a ticket costs $5.00, what is the expected profit per ticket?

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3820)


Part (a)

m = 1 first place prize
n = 1 second place prize
p = 3 third place prizes
m+n+p = 1+1+3 = 5 prizes total

This is out of 1,000,000 (one million) tickets sold.

The probability of winning a prize is 5/(1,000,000) = 1/(200,000) or "1 out of 200,000".

Answer as a fraction: 1/200000
Answer as a decimal: 0.000005

=====================================================================================

Part (b)

ScenarioProbability as a fractionProbability as a decimal
Win 1st prize1/(1,000,000)0.000001
Win 2nd prize1/(1,000,000)0.000001
Win 3rd prize3/(1,000,000)0.000003
Don't win anything(999,995)/(1,000,000)0.999995
The decimal values are exact without any rounding done to them.

X = net winnings aka profit

If a ticket costs $5, then we have these possible profits
$2,000,000 - $5 = $1,999,995
$250,000 - $5 = $249,995
$5,000 - $5 = $4,995
0-5 = -5 dollars

List each X with its corresponding P(X) value. I'll go with the decimal form of P(X).
XP(X)
1,999,9950.000001
249,9950.000001
49950.000003
-50.999995


Multiply each X value with its corresponding probability to form the X*P(X) column.
XP(X)X*P(X)
19999950.0000011.999995
2499950.0000010.249995
49950.0000030.014985
-50.999995-4.999975
Add up the items in the X*P(X) column to get
1.999995 + 0.249995 + 0.014985 + (-4.999975) = -2.735
This rounds to -2.74

With reference to the person who bought the lottery ticket, their expected profit per ticket is -2.74 dollars. They expect to lose, on average, about $2.74 per ticket.

Answer: -2.74 dollars


Answer by ikleyn(52943)

.
A lottery has a $ 2 000 000 first prize, a $ 250 000 second prize, and a three $5000 third prizes.
A total of 1 000 000 tickets are sold.
a) What is the probability of winning a prize in this lottery?
b) If a ticket costs $5.00, what is the expected profit per ticket?
~~~~~~~~~~~~~~~~~


As you formulate your problem,  I see that you are a beginner in composing  Math problems.

I say it because question  (a)  in the problem is posed incorrectly.

The correct formulation is

    +------------------------------------------------------+
    |   (a)  What is the probability of winning a prize    |
    |       in this lottery, having only one ticket ?      |
    +------------------------------------------------------+